OpenStudy (kailee1423):

The line y = 8x – 4 intersects the parabola y = x2 + 2x + 5 at exactly one point. What are the coordinates of the point of intersection?

OpenStudy (trojanpoem):

The intersection means that there is a point on both of the line and the parabola see the approximate figure: 

OpenStudy (trojanpoem):

So we can solve both equation together : 8x - 4 = x^2 + 2x + 5 x^2 - 6x + 9 = 0 (x - 3)(x-3) = 0 x = 3 y = 8 * 3 - 4 = 24 - 4= 20 So the point is (3,20)

OpenStudy (kailee1423):

Thank you,!!!!!

OpenStudy (trojanpoem):

You're welcome.